Standard Chartered Bank Zambia PLC (100late the earnings per share (EPS) as at 31st December 2019 and 2020 correct to two decimal places for Stanchart.
[10 Marks]
Calculate the earnings per share (EPS) as at 31st December 2019 and 2020 correct to two decimal places for the comparative company.
[10 Marks]
Explain Stanchart’s performance in 2020.
[10 Marks]

Explain fie (5) reasons for using EPS as a performance measure.
[10 Marks]
Explain 5 limitations of using ratio analysis as a means of assessing the bank’s performance.
[10 Marks]
1
Expert's answer
2021-09-10T00:19:55-0400

a.). EPS = "\frac{Net\; Income}{Outstanding\; Shares}"

EPS 2019:

Net income 2019 = 330,000

Income tax 2019 = 30"\% \times" 330,000 = 99,000

Net Income after tax = 330,000 – 99,000 = 231,000

Total outstanding shares 2019 = 100,000 shares


EPS 2019 = "\frac{231,000}{100,000} = 2.31"


EPS 2019 = K2.31

 

EPS 2020:

Net income 2020 = 250,000

Income tax 2019 = 30"\% \times" 250,000 = 75,000

Net Income after tax = 250,000 – 75,000 = 175,000

Total outstanding shares = 100000 + 50000 = 150,000 shares


EPS 2019 = "\frac{175,000}{150,000} = 1.17"


EPS 2020 = K1.17

 

b.). EPS for the comparative company: Zambia National Commercial Bank plc.

EPS 2019:

Net income 2019 after tax = 200,056,000

Total outstanding shares 2019 = 1,443,750,000 shares


EPS 2019 = "\frac{200,056,000}{1,443,750,000} = 0.14"


EPS 2019 = 0.14

 

EPS 2020:

Net income 2020 after tax = 206,658,000

Total outstanding shares 2020 = 1,443,750,000 shares


EPS 2020 = "\frac{206,658,000}{1,443,750,000} = 0.14"


EPS 2020 = K 0.14


(c). Explain Stanchart’s performance in 2020

- As comparing the Stanchart's performance in 2020 with performance in 2019 it is better because the EPS of the company is higher even after issuance of new shares in 2020.


(d). ) 5 reasons for using EPS as a performance measure.

  1. It measure profitability level. Earnings per share is a critical indicator for determining a company's profitability. In fact, EPS is the only statistic in the whole field of fundamental analysis that separates net income to determine what the shareholders earn by investing in the firm.
  2. It indicates the dividend of Payout. Dividends are a percentage of a company's profits that are paid out to shareholders. Many investors choose dividends because of the consistent income they provide. Dividends are also seen by investors as a favorable indicator of the company's future growth.
  3. It is a determines the P/E ratio. Aside from profitability, investors want to know how much value a specific stock adds to their portfolio. The P/E ratio can assist investors in determining the value of their stock holdings. EPS is the most important factor in determining the P/E ratio.
  4. It has high dimensional stability. The higher the EPS, the better the company's performance and prospects. The business's growth rate is reflected in the EPS track record over multiple years, and potential investors look forward to investing in the company if they observe an upward trend.
  5. It has a flexible structure. A weighted average of equity shares is determined if the number of equity shares fluctuates during the financial quarter. The number of days or months outstanding during the year determines the weight. Let's look at an example to assist us comprehend.


(e). 5 limitations of using ratio analysis as a means of assessing the bank’s performance.

  1. To enhance its ratios, the company can make certain year-end modifications to their financial statements. The ratios then become nothing more than window decoration.
  2. Inflationary pricing fluctuations are ignored by ratios. Many ratios are derived based on historical costs, and they ignore price changes between periods. This does not accurately reflect the financial position.
  3. Accounting ratios totally disregard the firm's qualitative elements. They solely evaluate the financial elements of the situation (quantitative).
  4. There are no universally accepted definitions for the ratios. As a result, various formulas for the ratios may be used by different companies. Current Ratio is one such example, where some companies count all current liabilities while others exclude bank overdrafts from current liabilities when calculating current ratio.
  5. Finally, accounting ratios do not address any of the company's financial difficulties. They are only a means to an aim, not a complete solution.